OpenStudy (anonymous):

Simplify the expression (properties of rational exponents) 4/(3)^-2

OpenStudy (anonymous):

is the ans is 28?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

or \[8\div3 ?\]

OpenStudy (anonymous):

the choices are 36x^2 or 24x^2

OpenStudy (lalaly):

your missing an x in ur question

OpenStudy (anonymous):

Use the formula: \[\frac{1}{b^{-a}} = b^a\]

OpenStudy (anonymous):

srry bout that, its where the 3 is.

OpenStudy (lalaly):

\[\frac{4}{(3x)^{-2}}\]is this it?

OpenStudy (anonymous):

So, the question is like this: \[\frac{4}{(3x)^{-2}}\]

OpenStudy (anonymous):

yes.

OpenStudy (lalaly):

\[\frac{1}{b^{-n}}=b^n\]soo\[\frac{4}{(3x)^{-2}}=4(3x)^2 =4 \times 9x^2 =36x^2\]

Parth (parthkohli):

\(\Rightarrow {4} \times \Large {1 \over (3x)^{-2}} \) \(\Rightarrow 4 \times (3x)^2 \) \(\Rightarrow 4 \times 9x^2\)

OpenStudy (anonymous):

thank u both, can you guys help me with another one?

Parth (parthkohli):

Ah sure.. close this one or post it right here

OpenStudy (anonymous):

1/5x times 1/5x

OpenStudy (anonymous):

I thought the ans was 1/25 but it's not so idk

OpenStudy (anonymous):

is it just 1/5x?

OpenStudy (lalaly):

\[\frac{1}{5x} \times \frac{1}{5x}=\frac{1}{5x \times 5x}=\frac{1}{25x^2}\]

Parth (parthkohli):

Nah.. \(\Large {1 \over 5x} \times {1 \over 5x} = {1 \times 1 \over 5x \times 5x}\)

Parth (parthkohli):

@lalaly you win at typing -_-

OpenStudy (lalaly):

lol parthy , next time i will let u do it

OpenStudy (anonymous):

so the ans would be 2/25?

OpenStudy (lalaly):

nno its 1/(25x^2)

OpenStudy (anonymous):

that's not one of my choices?

OpenStudy (anonymous):

I'll show u them

OpenStudy (anonymous):

1/5x , 2/10x, or 2/5x

OpenStudy (anonymous):

So, I am thinking you have to add that two numbers and not to multiply them.. \[\frac{1}{5x} + \frac{1}{5x} = \frac{1 +1}{5x} = \frac{2}{5x}\]

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

Welcome dear..